OpenStudy (anonymous):

Solve for x. 2-4/x^2=-2/x A. x = -2 B. x = 1 C. x = -2 or x = 1 D. x = 2 or x = -1

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

Answer is C :)

jimthompson5910 (jim_thompson5910):

the equation is \[\Large 2 - \frac{4}{x^2} = -\frac{2}{x}\] right?

OpenStudy (anonymous):

yep

jimthompson5910 (jim_thompson5910):

what did you get?

OpenStudy (anonymous):

I got C too, would that be correct?

jimthompson5910 (jim_thompson5910):

yes, C is correct, here's confirmation http://www.wolframalpha.com/input/?i=2+-+ \frac{4}{x^2}+%3D+-\frac{2}{x}

jimthompson5910 (jim_thompson5910):

clicking the link doesn't work, but copying/pasting does